# Rethinking cluster-conditioned diffusion models

### v1

@@ 5 Discussion
Figure 7: Generated low- (a) and high-confident (b) CIFAR10 samples . The top row depicts the unconditional (Uncond.) samples, while the bottom row shows the generated samples using C-EDM with TEMI ($C=100$). Images on the same column are produced with the same initial noise. Confidence is quantified using maximum softmax probability (MSP). MSP is measured using TEMI trained on CIFAR10 without annotated data.
Measuring the confidence of the generated samples using TEMI. In Fig. 7, we show the generated examples with the lowest and highest maximum softmax probability (MSP [33]) of the TEMI classification head as a measure of confidence. For comparison, we show unconditional samples that were generated using the same initial noise in the denoising process. Visual inspection shows that low-confidence C-EDM samples do not have coherent semantics compared to the unconditional ones, leading to inferior image quality. We hypothesize that the sampled condition for the low-confidence C-EDM samples is in conflict with the existing patterns in the initial noise. The above condition-noise conflict needs further investigation. By contrast, highly confident C-EDM samples show more clearly defined semantics than unconditional ones. When the low frequencies, such as the object’s shape, remain intact, cluster conditioning aids in refining local pixel patterns. Finally, we observe that highly confident C-EDM samples consist of simple pixel patterns that are easy to generate, such as a white background. More samples are provided in the supplementary material. To conclude, confidence can be leveraged in future works in various ways, such as rejection/acceptance sampling [20] or internal guidance methods [36, 37].
What about a lower cluster bound? Starting with a high overestimation of the number of clusters (e.g. $1K$ for CIFAR10), we find that TEMI clustering with $\gamma=1$ utilizes a subset of clusters, which could be used as a lower cluster bound. More precisely, we find a maximum standard deviation of $6.4$ for $C^{u}$ ($\gamma=1$) for multiple cluster values in the range of $[200,10K]$ across datasets and backbones (see Supp.). Intuitively, $C^{u}$ is the minimum amount of clusters TEMI (with $\gamma=1$) uses to group all image pairs while remaining discriminative. This behavior is analogous to cluster-based self-supervised learning (using image augmentations) [95, 18] and has been recently coined as partial prototype collapse [4]. Nonetheless, we consider the lower bound more applicable to large scales as the measured standard deviation might exclude the optimal granularity for small, highly curated datasets. Due to the above limitation and since the lower bound of $C$=2 adds minimal overhead using binary search (clustering-only iterations), we leave this for future work.
Image synthesis beyond ImageNet. ImageNet is currently the largest labeled public dataset, and a single experiment using a recent state-of-the-art diffusion model on ImageNet requires up to 4MWh at $512^{2}$ resolution [46]. Based on our experiments, clusters match or outperform the human-derived labels on image generation by estimating the visual groups. Using the introduced upper bound, the search space of the visual groups is significantly reduced with minimal computational overhead, while no further hyperparameter tuning is required. Therefore, it allows future works to incorporate unlabelled data and experiment at scales beyond ImageNet while being sample efficient. Additionally, the sample efficiency compared to noisy or non-mutually exclusive labels could be investigated in future works.
6 Conclusion
In this paper, a systematic empirical study was conducted focusing on conditioning diffusion models with cluster assignments. Given the optimal cluster granularity with respect to image synthesis (visual groups), it was demonstrated that cluster conditioning achieves state-of-the-art FID on three established generative benchmarks while attaining a strong sample efficiency. To reduce the search space for estimating the visual groups of the dataset, a novel method that computes an upper cluster bound based solely on clustering was proposed. Unlike the visual groups, which were found to be dependent on the generative metric, the discovered upper bounds were independent of the chosen metric. Finally, our experimental study points to the fact that there is no significant connection between the task of clustering and image synthesis.
Appendix 0.A Additional discussion points
0.A.0.1 Highly fine-grained clusters lead to out-of-distribution samples for classification datasets.
To provide further insights into the sensitivity of $C$ observed in Fig. 6, we report the AUROC using the top-1 NN cosine similarity [84] of the generated images against the official test split of CIFAR10 and CIFAR100 using DINO ViT-B (Fig. 8). The top-1 similarity is a metric of unsupervised out-of-distribution (OOD) detection [84]. Since the top-1 similarity is independent of the data distribution of the generated samples, it is independent of the diversity. In parallel, we measure how similar the cluster-conditional samples are compared to the unconditional ones using FID. We call this metric unconditional FID (uFID). A uFID of 0 would mean that the conditional and unconditional samples are identical w.r.t FID. Based on Fig. 8, we observe that uFID is increasing as $C$ increases, suggesting that the generated samples for larger $C$ are farther away in feature space than the unconditional samples. In parallel, the highest number of clusters produces the highest AUROC compared to the test split, suggesting that the features of the generated samples are recognized as OOD samples. We hypothesize that this behavior originates from highly specialized clusters that cannot always be generated from the initial noise, which results in visual artifacts. We leave further investigation on this for future work.
Figure 8: Top 1-NN similarity AUROC (left y-axis) and mean similarity score (MSS, right y-axis) w.r.t. the unconditional samples for different cluster sizes (x-axis) using C-EDM evaluated at $M_{img}=100$. Both metrics use DINO features.
0.A.0.2 FID and FDD across training iterations
@@
In Fig. 9, we report FID and FDD across training using C-EDM with TEMI clusters. We notice that FID tends to saturate faster than FDD and fluctuates more between checkpoints. FDD keeps decreasing monotonically, with very little fluctuation, and always prefers the samples at $M_{img}=200$. Since both metrics compute the Frechet distance, these tendencies can only be attributed to the supervised InceptionV3 features. Even though the study of generative metrics is out of the scope of this work and a human evaluation is necessary as in [83], we hope that our findings w.r.t. cluster-conditioning can facilitate future works.
@@
0.A.0.3 Dependence on $q(c)$ during generative sampling on balanced classification datasets.
EDM [44] Sampling Distribution CIFAR10 CIFAR100
$k$-means $\mathcal{U}(\{1,..,C\})$ $2.75$ $2.60$
$k$-means $q(c)$ $1.69$ $2.21$
$\Delta$ ($\downarrow$) - $0.79$ $0.39$
TEMI $\mathcal{U}(\{1,..,C\})$ $1.86$ $2.41$
TEMI $q(c)$ $1.67$ $2.17$
$\Delta$ ($\downarrow$) - 0.19 0.24
Table 4: We report FID for k-means and TEMI with and without taking into account the cluster distribution $q(c)$ of the training data. $\mathcal{U}(\{1,..,C\})$ denotes the uniform cluster distribution. We use $C_{V}=100,200,400$ for CIFAR10, CIFAR100 and FFHQ, respectively. $\Delta$ quantifies the absolute difference.
Appendix 0.B Deep image clustering with TEMI
0.B.1 Intuition for $\gamma$
@@
0.B.2 What about the lower bound? TEMI with $\gamma=1$ experiments.
@@
TEMI CIFAR10 FFHQ CIFAR100 $\gamma=1$ $C^{u}$ $C^{u}$ $C^{u}$ 100 33 36 48 400 38 48 48 500 34 54 51 800 40 45 47 1K 34 49 47 2K 48 52 54 5K 28 42 51 Mean 36.4 46.6 49.4 Std 6.37 6.16 2.63
@@
0.B.3 TEMI with different backbones.
@@
Appendix 0.C The EDM diffusion baseline.
@@
Appendix 0.D Additional implementation details and hyperparameters
@@
Table 7: Hyperparameters used for training EDM and CEDM. Bold signifies that the value is changing across datasets. All other parameters of the training setup were identical to the specifications of Karras et. al [44], which are detailed there.
@@
Appendix 0.E Visualizations
(a) High-confident C-EDM samples
(b) Low-confident C-EDM samples

### v2

@@ Appendix A Visualizations
Appendix B Additional discussion points
B.1 FID and FDD across training iterations
@@
In Fig. 15, we report FID and FDD across training using C-EDM with TEMI clusters. We notice that FID tends to saturate faster than FDD and fluctuates more between checkpoints. FDD keeps decreasing monotonically, with minimal fluctuation and always prefers the samples at $M_{img}=200$. Since both metrics compute the Frechet distance, these tendencies can only be attributed to the supervised InceptionV3 features. Even though the study of generative metrics is out of the scope of this work and a human evaluation is necessary as in [71], we hope that our findings w.r.t. cluster-conditioning can facilitate future works.
@@
B.2 Image synthesis beyond ImageNet.
ImageNet is currently the largest labeled public dataset, and a single experiment using a recent state-of-the-art diffusion model on ImageNet requires up to 4MWh at $512^{2}$ resolution [40]. Based on our experiments, clusters match or outperform the human-derived labels on image generation by estimating the visual groups. Using the introduced upper bound, the search space of the visual groups is significantly reduced with minimal computational overhead, while no further hyperparameter tuning is required. Therefore, it allows future works to incorporate unlabelled data and experiment at scales beyond ImageNet while being sample efficient. Additionally, the sample efficiency compared to noisy or non-mutually exclusive labels could be investigated in future works.
Appendix C Deep image clustering with TEMI
C.1 Intuition for $\gamma$
@@
C.2 What about the lower bound? TEMI with $\gamma=1$ experiments.
Starting with a high overestimation of the number of clusters (e.g. $1K$ for CIFAR10), we find that TEMI clustering with $\gamma=1$ utilizes a subset of clusters, which could be used as a lower cluster bound. More precisely, we find a maximum standard deviation of $6.4$ for $C^{u}$ across datasets and feature extractors (see Supp.). Intuitively, $C^{u}$ is the minimum amount of clusters TEMI (with $\gamma=1$) uses to group all image pairs. This behavior is analogous to cluster-based self-supervised learning (using image augmentations) [82, 15] and has been recently coined as partial prototype collapse [26]. Nonetheless, the lower bound is more applicable to large scales as the measured standard deviation might exclude the optimal granularity for small, highly curated datasets. Due to the above limitation, we leave this for future work.
@@
TEMI CIFAR10 FFHQ CIFAR100
$\gamma=1$ $C^{u}$ $C^{u}$ $C^{u}$
100 33 36 48
400 38 48 48
500 34 54 51
800 40 45 47
1K 34 49 47
2K 48 52 54
5K 28 42 51
Mean 36.4 46.6 49.4
Std 6.37 6.16 2.63
@@
C.3 TEMI with different backbones.
@@
C.4 Dependence on $q(c)$ during generative sampling on balanced classification datasets.
Table 6: We report FID for k-means and TEMI with and without considering the training data’s cluster distribution $q(c)$. $\mathcal{U}(\{1,..,C\})$ denotes the uniform cluster distribution. We use $C_{V}=100,200,400$ for CIFAR10, CIFAR100 and FFHQ, respectively. $\Delta$ quantifies the absolute difference.
EDM [39] Sampling Distribution CIFAR10 CIFAR100
$k$-means $\mathcal{U}(\{1,..,C\})$ $2.75$ $2.60$
$k$-means $q(c)$ $1.69$ $2.21$
$\Delta$ ($\downarrow$) - $0.79$ $0.39$
TEMI $\mathcal{U}(\{1,..,C\})$ $1.86$ $2.41$
TEMI $q(c)$ $1.67$ $2.17$
$\Delta$ ($\downarrow$) - 0.19 0.24
Appendix D The EDM diffusion baseline.
@@
Appendix E Additional implementation details and hyperparameters
@@
Table 7: Hyperparameters used for training EDM and C-EDM. Bold signifies that the value is changing across datasets. All other parameters of the training setup were identical to the specifications of Karras et. al [39], which are detailed there.
@@
Table 8: TEMI hyperparameters